OpenStudy (anonymous):

i need refreshing... can someone show me the steps for ------> √169? =13

OpenStudy (anonymous):

5

HanAkoSolo (jamierox4ev3r):

thats right? now what is \(5^2\)?

OpenStudy (anonymous):

25

OpenStudy (zzr0ck3r):

its the fact that 13*13 =168

OpenStudy (anonymous):

$$13^2=(10+3)13=130+3\times 13=130+39=169$$

OpenStudy (zzr0ck3r):

169*

HanAkoSolo (jamierox4ev3r):

right so if the square root of 25 is 5, and 5 squared is 25, then 13 squared is 169 and the square root of 169 is 13

HanAkoSolo (jamierox4ev3r):

you guys I was trying to solve here -.- and i was giving a verbal process

OpenStudy (zzr0ck3r):

2*2=4 sqrt(4) = 2 3*3 = 9 sqrt(9) = 3 4*4=16 sqrt(16) = 4 . . .

OpenStudy (anonymous):

hahaha calm down

HanAkoSolo (jamierox4ev3r):

lol :P jk

HanAkoSolo (jamierox4ev3r):

so \(13^2\) =169 and the square root of 169 also equals 13. Does that make sense now @panda03 ?

OpenStudy (anonymous):

yup ^_^ thanks everyone! especially @Jamierox4ev3r haha for the verbal process :D

OpenStudy (zzr0ck3r):

lol

HanAkoSolo (jamierox4ev3r):

lol yes my verbal process XD

OpenStudy (nincompoop):

OMG THIS QUESTION IS so hard :( I will practice more so I will be expert at it

OpenStudy (anonymous):

lol

HanAkoSolo (jamierox4ev3r):

uh oh here's Lord Nin.... XD

HanAkoSolo (jamierox4ev3r):

also @panda03 when in doubt, pull out the almighty calculator! lol

OpenStudy (anonymous):

i know lol, im practicing without the help of the calculator XD sooo yeaaa i said i need to refresh guys XD

HanAkoSolo (jamierox4ev3r):

lol well it is handy in a pinch :P

OpenStudy (anonymous):

haha yup thanks for your time tho :D